OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

An equation parallel to y=7x+2 would mean that they have the same slope. In slope-intercept form of y=mx+b, m is your slope. We now know that the new equation will also have a slope of 7. Afterwards you can plug in the point (1,-3) as values of the equation y=7x+b to find out b, which is the y- intercept. -3=7(1)+b -3=7+b b=-10 The final answer is A. 7X-10 (This is my first response, I hope it helps)

OpenStudy (anonymous):

Thank you And it did help (: